How much water must be mixed with 7 pints of 50% developer to make a mixture that is 13% developer?
Answer by ptaylor(2198)

Let x=amount of water needed
Now we know that the amount of pure developer in the water that's added (0) plus the amount of pure developer in the 7 pints of 50% developer (0.50*7) has to equal the amount of pure developer in the final mixture(0.13(7+x)). So our equation to solve is:
0+0.50*7=0.13(7+x) get rid of parens
0.50*7=0.91+0.13x subtract 0.91 from each side
3.5-0.91=0.13x
2.59=0.13x divide each side by 0.13
x=19.92 pints----amount of water needed
CK
0.50*7=0.13*(26.92)
3.5~~~~3.5
Hope this helps---ptaylor